In the laboratory a "coffee cup" calorimeter, or constant pressure calorimeter, is frequently used to determine the specific heat of a solid, or to measure the energy of a solutiorn phase reaction. Thermometer Stirring rod A student heats 63.92 grams of iron to 98.03 °C and then drops it into a cup containing 75.92 grams of water at 24.47 °C. She measures the final temperature to be 30.66 °C The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.57 J/°C. Water Metal Assuming that no heat is lost to the surroundings calculate thesample specific heat of iron Specific Heat (Fe) Submit Answer Retry Entire Group 4 more group attempts remaining
In the laboratory a "coffee cup" calorimeter, or constant pressure calorimeter, is frequently used to determine the specific heat of a solid, or to measure the energy of a solution phase reaction. Stirring rod Since the cup itself can absorb energy, a separate experiment is needed to determine the heat capacity of the calorimeter. This is known as calibrating the calorimeter and the value determined is called the calorimeter constant. One way to do this is to use a common metal of known heat capacity. In the laboratory a student heats 95.50 grams of zinc to 98.11 °C and then drops it into a cup containing 81.69 grams of water at 23.33 °C. She measures the final temperature to be 30.61 °C. ater Metal sample 2003 Using the accepted value for the specific heat of zinc (See the References tool), calculate the calorimeter constant. Calorimeter Constant- J/°C. Submit Answer Retry Entire Group 4 more group attempts remaining
In the laboratory a "coffee cup" calorimeter, or constant pressure calorimeter, is frequently used to determine the specific heat of a solid, or to measure the energy of a solutiorn phase reaction A chunk of lead weighing 19.83 grams and originally at 97.38 °C is dropped into an insulated cup containing 84.87 grams of water at 23.83°C. Stirring rod The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.78 J/°C. Metal sample Using the accepted value for the specific heat of lead (See the References tool), calculate the final temperature of the water. Assume that no heat is lost to the surroundings. final
When a solid dissolves in water, heat may be evolved or absorbed. The heat of dissolution (dissolving) can be determined using a coffee cup calorimeter. Thermometer In the laboratory a general chemistry student finds that when 4.06 g of NH4CI(s) are dissolved in 110.40 g of water, the temperature of the solution drops from 23.76 to 21.16 °C. Cardboard or Styrofoam lid The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.67 J/oC. Nested Styrofoam cup Based on the student's observation, calculate the enthalpy of dissolution of NH4Cl(s) in kJ/mol Reaction occurs in solution. Assume the specific heat of the solution is equal to the specific heat of water. kJ/mol dissolution
When a solid dissolves in water, heat may be evolved or absorbed. The heat of dissolution (dissolving) can be determined using a coffee cup calorimeter. Thermometer In the laboratory a general chemistry student finds that when 8.40 g of K2SO4(s) are dissolved in 102.30 g of water, the temperature of the solution drops from 24.96 to 21.94 °C. Cardboard or Styrofoam lid The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.59 Nested Styrofoam cups Based on the student's observation, calculate the enthalpy of dissolution of K,S04(s) in kJ/mol. Reaction occurs in solution. Assume the specific heat of the solution is equal to the specific heat of water. AH kJ/mol dissolution
When a solid dissolves in water, heat may be evolved or absorbed. The heat of dissolution (dissolving) can be determined using a coffee cup calorimeter. Thermometer In the laboratory a general chemistry student finds that when 0.87 g of KOH(s) are dissolved in 118.30 g of water, the temperature of the solution increases from 24.56 to 26.44 °C. Cardboard or Styrofoam lid The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.88 J/oC. Nested Styrofoam cup Based on the student's observation, calculate the enthalpy of dissolution of KOH(s) in kJ/mol Reaction occurs in solution. Assume the specific heat of the solution is equal to the specific heat of water. AHdissolution kJ/mol
Answer #1

1.SPECIFIC HEAT

S= q/m × ∆t

S= 1.57/ 63.92×98.03

S= 0.0025 j/ g° C

2. CALORIMETER CONSTANT

Energy lost by the zinc:

q = m Cp ΔT

q = (95.50 g) (0.39 J/g-1 °C-1) (98.11°C)

q = 3654.20 J

Energy gained by the water on adding zinc:

q = m Cp ΔT

q = (81.69 g) (0.39 J/g-1 °C-1) (23.33 °C)

q = 743.27 J

The calorimeter got the rest:

3654.20 - 743.27= 2920.93 J

Calorimeter constant :

2920.93 J / 30.61 °C =95.09 J/°C

3. FINAL TEMPERATURE

   We use: "q = mcΔT" except since assume q lead= qwater, you are simply solving for T, which is the final temperature. You need to look up the specific heat values (c) for lead and water. I used 0.160 for lead and 4.18 for water.

(19.83)(97.38 - T)(0.160) = (84.87)(T - 23.83)(4.18)

T = 24.51 °C

4. ENTHALPY OF DISSOLUTION for (NH4Cl)

Q= mc∆t

Q= 114.46 × 4.18 × (-2.16)

Q= -1033.43

Now ∆H = -Q/ Number of moles of solute

∆H= -(-1033.43/0.07) = 14763J

Or ∆H = 14.76 KJ

5. ENTHALPY OF DISSOLUTION for ( K2SO4)

Q= mc∆t

Q= 110.70 × 4.18 × (-3.02)

Q= -1397.43

Now ∆ H = -(-1397.43/0.048) = 29113J

Or ∆H= 29.11 KJ

6. ENTHALPY OF DISSOLUTION. for (KOH)

Q= mc∆t

Q= 119.17 × 4.18 × 1.88

Q= 936.48

Now ∆H= -(936.48/0.015)

∆H= -62432J = -62.43 KJ